### v2

Abstract
RNA function is deeply tied to secondary structure, with most molecules operating through dynamic and heterogeneous structural ensembles. While current analysis tools typically output single static structures or averaged contact maps, chemical probing methods like DMS capture nucleotide-resolution signals that represent the full structural ensemble, which however remain of difficult structural interpretation. To address this, we present MERGE-RNA, a framework that explicitly describes and outputs RNA as a structural ensemble. By modeling the experimental pipeline through a statistical physics framework, MERGE-RNA learns a small set of transferable and interpretable parameters, enabling the seamless integration of measurements across different molecules, probe concentrations, and replicates in a single optimization to improve robustness. Our model employs a maximum-entropy principle to predict thermodynamic populations, introducing only the minimal sequence-specific adjustments necessary to align the ensemble with experimental data. We validate MERGE-RNA on diverse RNAs, showing that it achieves strong structural accuracy and the ability to fill knowledge gaps in single-conformation reference structures. Furthermore, in a designed RNA construct for which we report new DMS data, MERGE-RNA deconvolves mixed states to reveal transient intermediate populations involved in strand displacement, dynamics that remain invisible to traditional analysis methods.
keywords:
RNA secondary structure, Chemical probing, Ensemble prediction, DMS-MaP
Graphical Abstract
RNA molecules play central roles in cellular processes, ranging from catalysis (ribozymes and ribosomes) [doudnaChemicalRepertoireNatural2002, ramakrishnanRibosomeStructureMechanism2002] to gene regulation (including riboswitches and microRNAs) [morrisRiseRegulatoryRNA2014]. This functional versatility arises from their ability to fold into specific secondary and tertiary structures [butcherMolecularInteractionsThat2011]. Importantly, RNA molecules exist as thermodynamic ensembles at both the secondary [boseCausesFunctionsTherapeutic2024] and tertiary [kenRNAConformationalPropensities2023] levels, with function arising from stable conformations, minority states, or dynamic equilibria. While tertiary structure modeling is computationally demanding [buRNAPuzzlesBlindPredictions2025], secondary structure thermodynamic models allow efficient computation of optimal structures [nussinovFastAlgorithmPredicting1980] and base-pairing probabilities [mccaskillEquilibriumPartitionFunction1990]. However, thermodynamic models face significant limitations from fundamental approximations: the nearest-neighbor approximation [tinocoImprovedEstimationSecondary1973], the neglect of pseudoknots [nussinovFastAlgorithmPredicting1980], and the hierarchical folding assumption [tinocoHowRNAFolds1999] may make it difficult to fully capture long-range interactions and context-dependent effects. To address these limitations, researchers have integrated co-evolutionary analysis or chemical probing data. Dimethyl sulfate (DMS) [zaugAnalysisStructureTetrahymena1995, kubotaProgressChallengesChemical2015] has become a powerful tool for structural analysis, providing single-nucleotide resolution data both in vitro and in vivo. DMS reacts preferentially with unpaired adenines and cytosines, with modifications detected as mutations during sequencing (mutational profiling, MaP) [kubotaProgressChallengesChemical2015]. However, interpreting this data is challenging: measured reactivities report on structural ensembles rather than single conformations, and experimental processes introduce systematic biases from reverse transcription and sequencing artifacts. Current approaches typically convert measured reactivities into pseudo-free-energy restraints [deiganAccurateSHAPEdirectedRNA2009], but this modifies predicted ensembles even when baseline models already agree with data [washietlRNAFoldingSoft2012], and the coefficients used to compute pseudo-free-energies are based on heuristics rather than physical principles.
Figure 1: Schematic overview of the method. Our approach builds a physical model—illustrated on the left side of the figure—that represents the entire pipeline of chemical probing experiments. This includes RNA folding into an ensemble of secondary structures, probe binding, adduct formation, and the final mutational profiling readout. The model is trained on chemical probing data from multiple RNA sequences, probe concentrations, and experimental replicates. From these data, it learns a set of physically meaningful parameters shared across all experiments ($\mu_{r}$, $\Delta\mu_{\text{pairing}}$, $p_{\mathrm{bind}}(A)$, $p_{\mathrm{bind}}(C)$, $p_{\mathrm{bind}}(G)$, $p_{\mathrm{bind}}(U)$, $m_{0}$ and $m_{1}$, as defined in the main text). In addition, it estimates sequence-specific soft constraints $\lambda_{i}$ (one per nucleotide position). Once trained, the model can predict the structural ensemble of each RNA at any probe concentration. Importantly, it can also extrapolate to zero probe concentration, providing the unperturbed structural ensemble.
Despite widespread practice, focusing on a single minimum free-energy (MFE) structure can be misleading as RNA molecules are known to adopt an ensemble of conformations, of which the MFE structure is often only a minor component. In fact, ignoring structural heterogeneity can pose problems both for interpreting RNA function and for properly modeling chemical probing experiments.
In this work we present MERGE-RNA (Multi-system Ensemble Refinement via Generalizable parameters Estimation), a unified physics-based model that describes chemical probing experiments with physically meaningful parameters and predicts RNA secondary structure ensembles. Our approach combines a standard thermodynamic folding model with explicit probe-RNA interactions characterized by two key parameters: a chemical potential ($\mu_{r}$) setting the probe concentration scale, and an energetic penalty ($\Delta\mu_{\text{pairing}}$) for probe binding to paired nucleotides. We employ maximum entropy inference to determine sequence-specific soft constraints ($\lambda_{i}$) representing minimal adjustments for model-data consistency, avoiding artificial constraints when baseline predictions already match data. The model explicitly accounts for mutational profiling artifacts, including false positives, false negatives, and position-specific biases. Crucially, our framework enables the simultaneous analysis of data from multiple RNA sequences, probe concentrations, and experimental replicates within a unified model, leveraging the shared physical parameters to improve robustness and extract transferable experimental knowledge.
In the next sections we outline the model and inference procedure, then validate transferability and accuracy on a group of well-characterized structured RNAs. We then apply the framework to the cspA 5 untranslated region (UTR) (a known case of temperature-dependent rearrangement). Furthermore, we demonstrate accurate deconvolution of mixed states in synthetic bistable constructs, and finally leverage experimental data to reveal intermediate, suboptimal populations involved in dynamic processes (e.g., strand displacement) that are missed by traditional approaches.
Methods
Physical Model
Our model follows the experimental pipeline of chemical probing experiments, as illustrated in Figure 1. The underlying physics remains consistent across experiments using the same reagent, probing time, and temperature, allowing us to simultaneously fit multiple experimental datasets within a unified framework. By maximizing the likelihood of the model to reproduce the observed mutation rates, our approach estimates physically meaningful parameters (detailed below) and identifies the ensemble of secondary structures most consistent with both the underlying physics and experimental evidence.
In this section we aim at introducing the physical model underlying our approach, while in the SI Appendix we provide a more in-depth description and rigorous discussion.
RNA folding and chemical probing
Folding
To model the RNA secondary structure ensemble, we utilize as baseline the ViennaRNA package [lorenzViennaRNAPackage202011]. The probability to observe a given secondary structure $s$ at temperature $T$ is defined as
$\mathbb{P}(s)=\frac{\exp\left(-\frac{F(s)}{\mathrm{k_{B}}T}\right)}{z}$ (1)
where $z$ is the partition function, $\mathrm{k_{B}}$ is the Boltzmann constant, and the baseline $F(s)=F_{0}(s)$ is obtained as a sum of thermodynamic parameters. To improve predictions, we refine the ensemble to align with experimental data while minimally perturbing the baseline model. Using the principle of maximum entropy [demartinoIntroductionMaximumEntropy2018], we optimize sequence-specific parameters ($\lambda_{i}$) that act as soft constraints for each nucleotide position [washietlRNAFoldingSoft2012]. These soft constraints modify the baseline ViennaRNA-predicted energy of pairing and remain consistent across experimental replicates and probe concentrations for a given RNA sequence. This results in a structure-dependent energy correction
$\Delta F_{\text{opt}}(s)=\sum_{i\in\text{paired}(s)}\lambda_{i}$ (2)
where $\text{paired}(s)$ denotes the list of paired nucleobases in structure $s$.
Probe-RNA interaction
We separately model the probe reactivity in two steps [calonaciMolecularDynamicsSimulations2023]: a reversible physical binding, described here, followed by an irreversible chemical reaction, as detailed in the next paragraph. At non-zero probe concentrations, we describe physical probe-RNA interactions through a thermodynamic binding model with two trainable parameters ($\mu_{r}$ and $\Delta\mu_{\text{pairing}}$). The reference chemical potential $\mu_{r}$ captures the probability of an unpaired nucleotide to bind to a chemical probe at a reference DMS concentration $\textrm{DMS}]_{r}=1\textrm{M}$. The chemical potential is then corrected based on the actual probe concentration [DMS] as $\mu=\mu_{r}+\mathrm{k_{B}T}\ln\left(\frac{[\text{DMS}]}{[\text{DMS}]_{r}}\right)$. $\Delta\mu_{\text{pairing}}$ differentiates between paired and unpaired sites, so that the effective chemical potential for paired nucleotides becomes $\mu^{\prime}\equiv\mu-\Delta\mu_{\text{pairing}}$. Since chemical probes typically exhibit reduced accessibility to base-paired regions, $\Delta\mu_{\text{pairing}}$ is expected to be positive, though the model could be straightforwardly generalized to cases where this is not true. Incorporating a penalty term for base-paired sites results in a modified partition function. This modification is implemented by applying a concentration-dependent energetic contribution to paired bases $\Delta F_{\text{[DMS]}}(s)$ (derivation is detailed in the SI Appendix).
Putting together all these terms, the corrected free energy of a given secondary structure $s$ is
$F(s)$ $=\underbrace{F_{0}(s)}_{\text{baseline}}+\underbrace{\sum_{i\in\text{paired}(s)}\lambda_{i}}_{\begin{subarray}{c}\text{site-specific}\\ \text{corrections}\end{subarray}}+\underbrace{\Delta F_{\text{[DMS]}}(s)}_{\begin{subarray}{c}\text{concentration}\\ \text{dependent perturbation}\end{subarray}}$ (3)
from which we can compute the population of observing each structure $s$ via Equation 1.
Chemical binding
Following the physical interaction between a chemical probe and RNA, adducts are formed with probabilities that depend on both nucleobase identity and local RNA structure, under fixed experimental conditions (e.g., temperature and probing time). In the most general framework, this yields eight possible configurations, defined by the combination of nucleotide type (A, U, C, or G) and structural state (paired or unpaired). For DMS probing, due to the position of the potentially reactive nitrogens, adenine and cytosine show negligible reactivity when base-paired, while guanine and uracil retain reduced but secondary structure-independent reactivity. Accordingly, the model is parameterized by four effective binding probabilities, corresponding to the reactive nucleotide-structure combinations: $p_{\text{bind}}(\text{A},\text{unpaired})$, $p_{\text{bind}}(\text{C},\text{unpaired})$, $p_{\text{bind}}(\text{G})$, and $p_{\text{bind}}(\text{U})$.
Mutational Profiling and prediction of reactivity profiles
Mutational profiling
Following chemical probing, modifications are typically detected through reverse transcription. Here, we focus on mutational profiling (MaP), though only minor modifications to the model would be required for compatibility with other experimental techniques, such as RT-stop. During reverse transcription, we account for the probability of producing false positive or false negative signals through the parameters $m_{0}$ and $m_{1}$, as detailed in the SI Appendix. Applying this framework to a given secondary structure $s$, we can compute the expected reactivity profile for that structure, then sum over all secondary structures in the ensemble to obtain the expected reactivity profile for the experiment.
Handling systematic errors
Sequencing errors are known to introduce systematic biases in reactivity profiles [liu-weiSequencingAccuracySystematic2024], and reverse transcription can be affected by the presence of secondary structures [verwiltArtifactsBiasesReverse2023]. To account for such effects, we make use of experimental controls (performed without reagent) to estimate position-dependent correction factors $\epsilon_{i}$, which effectively act as a baseline adjustment to the predicted mutation rates at each nucleotide position and are computed so that the prediction at zero concentration of DMS matches the control data. These factors capture systematic biases that are sequence-specific but independent of probe concentration. This allows the model to focus on the concentration-dependent signal in reactivity, which is most informative about RNA structure and probe interactions, while $\epsilon_{i}$ absorbs background effects that remain constant across conditions. In other words, the sequence-dependent $\epsilon_{i}$ parameters model the mutation rate in absence of reagents, whereas the sequence-dependent soft constraints $\lambda_{i}$ affect the pairing population which, in turn, controls how much the mutation rate increases when the reagent is added. As discussed in the SI Appendix, when control experiments are available the parameter $m_{0}$ is omitted, since background correction factors can be directly estimated from the control data. If controls are unavailable, $m_{0}$ is instead fitted to provide a single background correction factor shared across the sequence.
Parameter Optimization
Upon computation of the expected reactivity profile for a given sequence and probe concentration, we optimize the model parameters by minimizing a loss function that quantifies the discrepancy between predicted and observed reactivities. We model per-position mutation counts as binomial: given read depth $n_{i}$ and model-predicted mutation probability $M_{i}$, each read is an independent Bernoulli trial. Accordingly,
$\mathbb{P}\big(\mathcal{M}_{i}^{\mathrm{exp}}\mid n_{i},M_{i}\big)$ $=\mathrm{Binomial}\big(\mathcal{M}_{i}^{\mathrm{exp}};n_{i},M_{i}\big)$ (4)
where $\mathcal{M}_{i}^{\mathrm{exp}}$ is the experimentally observed number of mutations at position $i$. As discussed in the SI Appendix, the overall likelihood of the model across all nucleotide positions is the product of the individual position likelihoods. To facilitate optimization, we minimize the negative log-likelihood, yielding the following loss function:
$\mathcal{L}$ $=-\sum_{i}\ln\,\mathbb{P}\big(\mathcal{M}_{i}^{\mathrm{exp}}\mid n_{i},M_{i}\big).$ (5)
The derivation and all definitions are expanded in the SI Appendix.
This optimization is implemented via a gradient-based algorithm that systematically adjusts model parameters to minimize the discrepancy between predicted and experimental reactivity profiles [byrdLimitedMemoryAlgorithm1995] as implemented in SciPy [virtanenSciPy10Fundamental2020]. The parameter optimization is performed in two consecutive phases, as illustrated in Figure 1. In the first phase, we optimize the global physical parameters that are shared across all sequences and concentrations. These include: the reference chemical potential ($\mu_{r}$), the binding penalty for structured regions ($\Delta\mu_{\text{pairing}}$), the nucleotide-specific adduct formation probabilities ($p_{\text{bind}}(\text{A},\text{unpaired})$, $p_{\text{bind}}(\text{C},\text{unpaired})$, $p_{\text{bind}}(\text{G})$, and $p_{\text{bind}}(\text{U})$), the false negative rate parameter ($m_{1}$), and—when experimental controls are unavailable—the background correction factor ($m_{0}$). This phase optimizes a total of 7 parameters (8 if no controls are available). In the second phase, we look to refine the secondary structure ensemble to find the one that best explains the experimental data. This is achieved by optimizing the sequence-specific soft constraints ($\lambda_{i}$) for each nucleotide position, while keeping the global physical parameters fixed.
Following parameter optimization, $\Delta F_{\text{[DMS]}}(s)$ is set to zero, extrapolating to zero probe concentration and thereby revealing the ensemble of secondary structures most consistent with the experimental data under native conditions.
Experimental procedures
DMS probing of a synthetic RNA construct was performed as described in the appendix. DMS-treated RNA was column-purified and subjected to reverse transcription using MarathonRT [zhaoUltraprocessiveAccurateReverse2018, guoSequencingStructureProbing2020, bohnNanoDMSMaPAllowsIsoformspecific2023, gribling-burrerIsoformspecificRNAStructure2024]. Sequencing reads were processed with an RNAFramework-based workflow [incarnatoRNAFrameworkAllinone2018, langmeadFastGappedreadAlignment2012, danecekTwelveYearsSAMtools2021].
Results
Transferability and accuracy of the physical model
We extracted physical parameters from our unified physics-based framework after training on a set of well-characterized RNA structures previously studied using DMS chemical probing [bohnNanoDMSMaPAllowsIsoformspecific2023]. This set spans a range of structural complexity, comprising five distinct systems: (i) bacterial RNase P (type A), (ii) hc16 ligase, (iii) Tetrahymena ribozyme, (iv) V. cholerae glycine riboswitch, and (v) HCV IRES. These sequences range from approximately 100 to 400 nucleotides in length. Reference structures for hc16 ligase, Tetrahymena ribozyme, and V. cholerae glycine riboswitch were obtained from cryo-EM-guided structure determination [kappelAcceleratedCryoEMguidedDetermination2020], and secondary structure was annotated from the 3D structure with Barnaba [bottaroBarnabaSoftwareAnalysis2019]. HCV IRES reference structure was manually transcribed from Extended Data Fig. 6 of Ref.  [bohnNanoDMSMaPAllowsIsoformspecific2023]. For bacterial RNase P (type A), although a crystal structure exists [torres-lariosCrystalStructureRNA2005], sequence differences between the crystallized construct and the chemically probed sequence preclude its use as a reference. These reference structures are expected to be more accurate with respect to the predictions done using a simple thermodynamic model. However, since they represent single structures, they do not report on conformational heterogeneity.
Figure 2: (a) Cross-validation of physical parameters across five RNA systems. Each bar represents the per-datapoint loss obtained for the indicated system after training on three systems and testing on the remaining two. Yellow bars refer to results obtained when the corresponding system was excluded from the training set. (b) Normalized loss profiles for HCV IRES with soft constraints of varying magnitude applied. Red line: constraints from reference structure; light green line: constraints from minimum free energy (MFE) structure. Each curve represents an independent fit with physical parameters trained on different system triplets. The blue line corresponds to an improved fit achieved by optimizing the sequence-specific $\lambda_{i}$ parameters with MERGE-RNA. Stars are reused across panels to indicate the same ensemble. (c) Frobenius distance divided by the squared number of nucleotides between base pairing probability (bpp) matrices, quantifying structural similarity between the HCV IRES reference structure and: MFE structure, bpp of the original thermodynamic ensemble, bpp obtained from the minimum of the red curve in panel (b), and bpp from the fully optimized model. (d) Mutation rates at 57mM after background subtraction for adenine and cytosine in HCV IRES, grouped by pairing status in the reference structure. Each point is colored according to the pairing probability predicted by MERGE-RNA. Strong disagreements (positions where MERGE-RNA predicts pairing probability below 25% or above 75% while the reference indicates the opposite) are highlighted.
Figure 3: Arc plot visualization comparing the secondary structures of HCV IRES: reference experimental structure (red), ensemble obtained from the thermodynamics model (green), and ensemble obtained from MERGE-RNA (blue).
Our first objective was to assess whether the physical parameters extracted from our model represent genuine physical properties of the experiment and are hence transferable to different systems, rather than being overfitted on a specific case. We performed comprehensive cross-validation fits, training the physical parameters of the model (all the parameters except the $\lambda_{i}$) on triplets of RNA systems and testing on the remaining pairs (Fig. 2a). Loss values (as defined in Equation 31) are then divided by the number of data points to enable direct comparison. The loss values obtained across all sets are consistent both in training and in test, demonstrating that our model generalizes on multiple systems and successfully captures the underlying principles governing chemical probing experiments. Notably, physical parameters obtained through different minimizations might be different (see Table S1) but typically predict similar mutation profiles. One intuitive example of this is that by increasing the reference chemical potential ($\mu_{r}$), the predicted mutation rate will–on average–increase, but this can be compensated by a decrease of binding probability $p_{\text{bind}}$. This under-determination of parameters can arise when the experiments do not cover a wide enough range of probe concentrations, but it does not affect the model’s ability to predict reactivity profiles within the range of conditions upon which it was trained.
Importance of conformational heterogeneity in structured RNAs
Having established the generalizability of the physical parameters, we next used our framework to explore the balance between structural accuracy and ensemble heterogeneity. Using HCV IRES as a representative example (analogous results for other systems in Supplementary Figs. S1,S2,S3), we varied a uniform soft constraint (Fig. 2b): for a given scalar magnitude (x-axis), every nucleotide annotated as paired in the chosen reference structure (experimental or minimum-free-energy) was assigned that value and all other positions were set to zero, with no position-specific tuning. Increasing this magnitude progressively biases the ensemble toward the targeted structure, smoothly interpolating between the original thermodynamic ensemble and a single constrained structure while keeping the physical parameters fixed.
By interpolating between the original ensemble and its single most-likely structure (minimum-free-energy), we assess the impact of reducing ensemble heterogeneity on the agreement with experiment (light green line in Fig. 2b). In fact, favoring a single structure has the effect of progressively suppressing the population of alternative conformations that are likely present in solution and contribute to the observed mutation profiles. Interestingly, the loss function monotonically increases when the constraint is applied, degrading the agreement with experiment. On the other hand, incorporating knowledge from the reference experimental structure (red line in Fig. 2b) initially improves the model’s agreement with experimental data. However, even in this case, excessive constraint enforcement eventually degrades performance. This pattern indicates the existence of a “sweet spot” where the model partially incorporates reference structural information while preserving the conformational diversity necessary to explain the experimental measurements. Too strong a constraint toward any single structure—even the experimental reference—reduces the model’s ability to account for the heterogeneous structural ensemble that is captured in the experimental data. Overall, this analysis shows that a proper balance of information from the reference single experimental structure and ensemble diversity improves the capability of the model to reproduce chemical probing experiments.
Inference of optimal ensemble
Building on these insights, instead of borrowing from the reference structure, we inferred the best sequence-specific soft constraints ($\lambda_{i}$ parameters) using our model. This resulted in substantially improved agreement with experimental data (blue line in Fig. 2b), as expected given that the experimental data are used to infer the soft constraints.
To validate our model independently, we computed the Frobenius distance between the reference structure and base-pairing probability matrices obtained by the baseline ViennaRNA MFE, Vienna BPP, and those optimally guided by the reference structure (Fig. 2c). Our model achieves comparable accuracy for some systems and clear improvements for others (Figs. S1–S3 in SI Appendix). Remarkably, our inference performs strongly even against ensembles obtained using uniform soft constraints that are reference-informed (red star on the red curve in Fig. 2b) yet our model does this without knowledge of the underlying experimental structure. This performance is likely due to our model fitting each nucleotide’s soft constraint ($\lambda_{i}$) individually, whereas the reference-informed approach applies a uniform constraint to all paired positions. Arc plots (Fig. 3) illustrate how this MERGE-RNA’s individualized fitting preserves ensemble diversity (blue) while recapitulating the reference structure (red) more accurately than the standard thermodynamic ensemble (green). Similar improvements were observed across all tested RNA systems, confirming the broad applicability of our approach.
MERGE-RNA reconciles dynamic probing data with static reference structures
Having established structural accuracy by comparison with experimentally resolved reference structures, we now examine where our predictions diverge from the static reference. Since our model predicts a dynamic ensemble rather than a single conformation, divergence is expected and informative: chemical probing experiments inherently report on the dynamic ensemble of the RNA molecule, which a static reference structure alone cannot fully capture. In Figure 2d we report mutation rates after control subtraction for adenine and cytosine, grouped by pairing status in the reference structure. As expected, paired positions show, on average, lower mutation rates than unpaired ones. Each point is colored according to the pairing status predicted by MERGE-RNA. The model’s predictions correlate well with mutation rates, demonstrating its capability to infer and model pairing dynamics from probing data. It is instructive to examine positions where our predictions disagree with the reference structure.
When mutation rates are consistent with the reference structure, MERGE-RNA’s predictions align with it. For positions with intermediate mutation rates, MERGE-RNA predicts intermediate pairing probabilities, suggesting conformational heterogeneity in solution. This is particularly informative: a static reference structure classifies each position as either paired or unpaired, and thus cannot account for intermediate reactivity. MERGE-RNA bridges this knowledge gap by proposing alternative conformations that together explain the observed mutation rates. Where probing data and reference structure are in contrast, MERGE-RNA captures new pairings absent from the reference but supported by probing, or identifies reference-annotated pairs that show elevated mutation rates, suggesting they are frequently unpaired in the dynamic ensemble.
The number of “strong disagreements sites” (positions where MERGE-RNA predicts pairing probability below 25% or above 75% while the reference indicates the opposite for A and C nucleotides, highlighted in Fig. 2d) is modest, ranging from 0 to 13 across the systems studied here. Detailed analysis (SI Appendix Figs. S4S7) reveals that most correspond to small corrections of existing structural elements or formation/dismantling of short motifs (e.g., 3–4 bp stems), indicating localized inconsistencies rather than wide structural disagreements. We note that some discrepancies may also arise from in vitro versus in-cell conditions, protein binding, errors in reference structure annotation, or other artifacts.
In summary, with our model (1) the inferred physical parameters are generalisable across systems (Fig. 2a); (2) we show that there is an optimal balance between structural accuracy and ensemble heterogeneity, and inference of sequence-specific corrections further improves agreement with experimental data (Fig. 2b); (3) the resulting structural ensembles more accurately capture native conformations compared to standard thermodynamic approaches (Fig. 2c); (4) comparing MERGE-RNA’s results to the reference structures reveals that agreement reflects probing data consistency (Fig. 2d). In cases where reference models are inconsistent with experiment, MERGE-RNA goes beyond the static reference by identifying localized discrepancies and proposing alternative conformations that better explain the data. (5) these results are consistent across all tested RNA systems (SI Appendix Figs. S1S3).
Application on a known case of functional structural rearrangement: cspA 5’ UTR
Figure 4: Ensemble predictions for cspA 5 UTR at two different temperatures. (a,b) Arc plots corresponding to the DMS data collected at 10 and 37C, respectively. Predictions from MERGE-RNA (red and black) are compared with predictions from Ref. [zhangStressResponseThat2018] (blue).
We next tested our model’s ability to capture complex structural rearrangements in relevant biological systems by considering two DMS mutational profiles for cspA 5 UTR recorded at 10 and 37C [zhangStressResponseThat2018]. This sequence acts as a thermoswitch [giuliodoriCspAMRNAThermosensor2010], with different structures at the two reported temperatures. The two experiments have been recorded at different probing times (2 hours vs 10 minutes). Our approach has been designed to analyze homogeneous experiments and does not model the dependence of the observed mutation rate on the probing time. We hence normalize the dataset at 10C to the same average mutation rate observed at 37C. In addition, to avoid confusing the effect of the mutational profiles with the temperature-dependence of the thermodynamic model, we use the thermodynamic model set on the same intermediate temperature for every ensemble (23.5C). Finally, we note that these experimental datasets lack a control mutation profile obtained in absence of DMS, making the separation of the physical signal from the systematic bias more difficult for our model. Given the mentioned limitations, we decided to perform a set of 6 independent fits, starting from different random initializations of the physical parameters. We report in Fig. 4 the ensemble predicted by the model that achieved the lowest loss, and the other fits are reported in the SI Appendix (Figs S8S12).
Figure 4(a,b) shows the structural ensembles predicted by MERGE-RNA, which largely recapitulate the individual structures reported by Zhang et al. [zhangStressResponseThat2018]. At 10C and 37C, our ensembles closely match their predictions in positions 1–120 and 70–170, respectively, but exhibit notable differences in other regions, particularly in segments where alternative helices compete. Unlike Ref. [zhangStressResponseThat2018], which reports a single minimum-free-energy structure, MERGE-RNA predicts a thermodynamic ensemble that accounts for multiple competitive helices and their relative populations. Overall, this analysis demonstrates that our model can capture complex structural rearrangements in biological systems.
Deconvolution of mixed structural states on synthetic data
Figure 5: Model accurately deconvolves mixed structural states from synthetic data. (a) Base-pairing probabilities inferred from fitting on synthetic data for a bistable RNA sequence. The data was generated from an RNA construct with two competing helices (helix 1 and helix 2), where the ground-truth population of helix 1 was systematically varied from 0 to 11 with steps of $0.2$. The model’s predicted pairing probabilities are shown for each case. (b) Arc plots of the inferred structural ensemble for two limit cases: when helix 1 is dominant (gold) and when helix 2 is dominant (black). (c) The helix population inferred by the model is plotted against the ground-truth population. The inferred population is quantified as the median of the base-pairing probabilities over the respective helical regions, highlighted in panels (a) and (b).
To rigorously test the model’s ability to correctly predict secondary-structure populations in heterogeneous ensembles, we generated synthetic datasets for an in-house designed bistable RNA carrying two mutually exclusive hairpins. We deliberately designed this construct so that we could also assay the very same sequence experimentally (discussed in the next section); accordingly, we appended primer binding sites at the 3 and 5 ends. Concretely, the final construct has the architecture
$5^{\prime}\text{--PBS1--}\text{A1--L1--B}\text{--L2--A2}\text{--PBS2--}3^{\prime}$ (6)
where PBS1 and PBS2 are distinct primer-binding sites, A1 and A2 are identical copies of the same sequence, and B is complementary to A1/A2. In this design, A1 and A2 compete to pair with B, producing two mutually exclusive hairpins (helix 1: A1:B; helix 2: A2:B; as schematically visualized in Fig. 6c, see ”baseline prediction”). L1 and L2 are short linkers designed to act as tetraloops for the hairpins.
While the original design targeted roughly equal populations of the two helices, the thermodynamic model allows us to artificially bias the ensemble to favor either helix. We then simulated mutation profiles for the two cases—when either helix is always formed—and combined these profiles in known proportions to create synthetic datasets with varying helix populations, ranging from 0 to $1$ in increments of $0.2$. To generate realistic mutation profiles, we used the physical parameters inferred from our analysis on experimental data from biologically relevant RNAs (subsection ’Transferability and accuracy of the physical model’ above).
We then trained our model on these synthetic datasets. Following the procedure established earlier, we first performed a joint fit of the physical parameters across all datasets, starting from random initializations. Subsequently, we optimized the soft constraints $\lambda_{i}$ for each dataset individually. The results, presented in Fig. 5, show that the model accurately recovers the expected base-pairing probabilities for both helices across all synthetic datasets. The method is robust to the presence of additional base pairs formed by the random primer binding sites. Overall, these synthetic data tests validate the model’s ability to resolve heterogeneous RNA ensembles and recover helix populations consistent with the known ground truth (Fig. 5c).
A putatively bistable sequence exhibits a heterogeneous ensemble with strand displacement
Figure 6: Ensemble inference on experimental data for a putative bistable RNA exhibits evidence of strand displacement. (a) Per-nucleotide DMS-MaP mutation profiles across 0–100 mM (20 mM steps; two replicates each) for a putative bistable RNA. We leave out of the plot the pairing binding sites for better visualization. Two regions (A1 on the left and A2 on the right, as defined in (6)) are competing to pair with the central region (B). They show a slope in the reactivity profile that is compatible with a strand displacement mechanism. (b) Pairing probabilities obtained from the ensemble predicted by MERGE-RNA (solid lines) after fitting on experimental data, with Turner 2004 [mathewsIncorporatingChemicalModification2004] (crimson) or Andronescu 2007 [andronescuEfficientParameterEstimation2007] (steel) thermodynamic models as baselines. MERGE-RNA is able to fix both the baseline predictions (dotted lines) and retrieve pairing probabilities consistent with the observed mutation rates. (c) Schematic depiction of the ensemble of structures as predicted by the baseline thermodynamic model (above) and MERGE-RNA fitted on the experimental data (below). While the thermodynamic model predicts only the two configurations with fully-formed helices, MERGE-RNA captures a wealth of intermediate states where strand displacement occurs.
To validate our model on experimental data, we collected DMS-MaP data on the putative bistable RNA construct described in the previous section. The data were collected at multiple DMS concentrations, 0 to 100 mM at steps of 20 mM, 2 replicas each. The obtained mutation profiles are visualized in Fig. 6a. We then trained our model on these experimental datasets, following the same procedure as in the previous section. Note that the sequence, as discussed earlier, was designed to adopt two mutually exclusive helices. We then applied our model to infer the structural ensembles from the data.
To challenge our model’s ability to recover the correct structural ensemble from an inaccurate thermodynamic baseline ($F_{0}$ term in Equation 3), we performed the fitting procedure using two different baseline models. We first used the default Turner 2004 parameters [mathewsIncorporatingChemicalModification2004], which were employed in the original sequence design. We then repeated the analysis with the Andronescu 2007 parameters [andronescuEfficientParameterEstimation2007], which predict a substantially different initial ensemble with populations of approximately 20% helix 1 and 80% helix 2, in stark contrast to the roughly 50-50 distribution predicted by the Turner 2004 parameters, as illustrated in Fig. 6b, dotted lines.
The results of our model fitting are presented in the same panel as solid lines. Remarkably, the two predictions of MERGE-RNA are consistent with each other, regardless of the initial thermodynamic parameter used, signaling robustness of the prediction with respect to the baseline parameters. In both cases, the model adjusts the ensemble of structures so that the resulting ensemble slightly favors helix 1, in accordance with the lower mutation rates observed in the experimental data (Fig. 6a).
Crucially, MERGE-RNA does more than adjust the populations of the two competing helices. It reveals a gradient in base-pairing probabilities within regions A1 and A2, where nucleotides closer to the central domain (B) are more likely to be paired, reflecting the slope observed in the experimental mutation profiles (Fig. 6a). Beyond the two fully-formed helices, the model resolves intermediate states in which B is partially paired with both A1 and A2, as schematically illustrated in Fig. 6c. This mechanism, known as strand displacement (or strand invasion) [ratajczykControllingDNARNA2025], has been observed in biological systems and is consistent with the observed mutation profile: regions nearest the central domain are expected to pair first during displacement, and accordingly show reduced reactivity closer to the loops (Fig. 6a). The presence of such intermediate states is expected if the two loops are sufficiently stable. Analysis of loop co-occupancy (Table S2) demonstrates capture of intermediate states missed by standard models. While baseline ensembles rarely allow both loops to form (<0.2<0.2%), the model predicts simultaneous pairing in 42–57% of structures, consistent with the observed mutation profiles and reflecting the strand-displacement mechanism in which the central domain transiently pairs with both competing helices. Overall, these results demonstrate MERGE-RNA’s ability to capture a continuous spectrum of low-population intermediates that standard thermodynamic models fail to predict, highlighting its ability to reveal dynamic structural ensembles inaccessible to conventional approaches, independently of the baseline thermodynamic model.
Discussion
In this work, we have introduced MERGE-RNA, a physics-based framework for the analysis of chemical probing data that addresses fundamental limitations of conventional approaches. Our approach moves beyond the use of experimental data as simple pseudo-energetic restraints [deiganAccurateSHAPEdirectedRNA2009] and instead models the entire experimental pipeline, from the physical interaction of the probe with the RNA to the final mutational readout. This enables the determination of transferable, physically meaningful parameters and the integration of multiple datasets for inferring the native structural ensemble. As demonstrated by our successful deconvolution of mixed structural states (both on synthetic and experimental data), our method provides a powerful tool for quantitatively characterizing conformational populations, crucial for understanding functional RNAs and dissecting processes such as strand displacement and co-transcriptional folding.
We note that our model explicitly includes all four nucleotides even though DMS exhibits structure-dependent reactivity primarily with A and C. Including the complete sequence leads to more robust estimates of the physical parameters, ultimately improving the accuracy of inferred structural ensembles.
A central strength is the ability to avoid over-constraining: standard pseudo-free energy approaches [deiganAccurateSHAPEdirectedRNA2009] modify the energy landscape globally, often altering already-correct regions. Similarly to Ref. [washietlRNAFoldingSoft2012], our framework infers minimal soft constraints, ensuring modifications arise only where data support deviation, and resulting in a lower loss (Fig. 2b). In addition, our physics-based model naturally enables the combination of multiple datasets in a single fitting process. Among ensembles compatible with experiment, the maximum entropy principle chooses the one closest to the prior, and modifications of the ensemble are here regularized by bounding soft constraints. Alternative regularization approaches might be explored [hummerBayesianEnsembleRefinement2015, cesariUsingMaximumEntropy2018].
Existing ensemble-deconvolution pipelines such as DREEM [tomezskoDeterminationRNAStructural2020], DRACO [morandiGenomescaleDeconvolutionRNA2021], and DANCE-MaP [olsonDiscoveryLargescaleCellstateresponsive2022] cluster sequencing reads with multiple co-occurring modifications and assign each cluster to a discrete structure. By capitalizing on per-read correlation patterns, these methods recover information about alternative conformations but at the cost of discarding the abundant reads with zero or single modifications and demanding deep coverage to resolve low-probability states. Alternative ensemble fitting approaches include Rsample [spasicModelingRNASecondary2018] and SLEQ [liStatisticalModelingRNA2018]. Rsample employs a sample-and-cluster approach, reweighting the Boltzmann ensemble through pseudo-free energy adjustments resembling our soft constraints, though without physics-grounded calibration, but rather adopting an empirical strategy in a non self-consistent fashion. SLEQ uses a sample-and-select strategy rooted in the maximum parsimony principle to identify the sparsest set of structures explaining modification patterns, differing from our maximum entropy approach designed to minimally perturb the thermodynamic model. Our framework calibrates against averaged mutational fractions, ensuring every read (even negatives) contributes signal in sparse-modification regimes. The trade-off is loss of explicit per-read co-variation, though reads could in principle be integrated to fine-tune populations—a possible future extension. Read-level clustering would struggle to distinguish multiple similar conformations and represent continuous ensembles with many intermediates (e.g., strand-displacement pathways). Furthermore, alternative methods assume each read originates from a single structure, which may not hold for RNAs undergoing rapid dynamics. Our framework models ensemble-averaged behavior directly, accommodating fast interconversions at equilibrium. Consequently, our framework captures both discrete alternative structures and continuous ensembles enabling interpretation of dynamic processes. Strikingly, even without individual reads, the model reconstructs correlations and measures minor substate populations, which would be missed in methods that enumerate a small number of representative structures.
A key feature is predicting the structural ensemble at any probe concentration, especially extrapolating to zero concentration for the native state. In our analyses, base-pairing probability dependence on concentration was modest, introducing partial parameter under-determination (e.g., increasing $\mu$ mimics decreasing $p_{\mathrm{bind}}$). This does not compromise robustness or transferability (Fig.2a) and can be mitigated by widening the concentration range; Bayesian treatment could be used to quantify identifiability. At the same time, the model is equipped to handle systems where probes induce significant non-linearities [calonaciMolecularDynamicsSimulations2023] from rearrangement, denaturation, or saturation. Standard protocols recommend avoiding these effects by reaching the single-hit-kinetics [arnoldCooperativityRNAChemical2023], which might be difficult or impossible for large RNAs, and intrinsically eliminates correlation information. Explicitly modeling these concentration-dependent perturbations might enable a more faithful representation of native conformational landscapes.
The separation of RNA thermodynamics (baseline model and $\lambda_{i}$) from probe-specific physics (all fo the other parameters) enables future multi-probe extensions. One could analyze data from CMCT [ehresmannProbingStructureRNAs1987], SHAPE [merinoRNAStructureAnalysis2005, weeksAdvancesRNAStructure2010, marinusNovelSHAPEReagent2021], hydroxyl radical footprinting [adilakshmiHydroxylRadicalFootprinting2006], and ETS [doudsNewReagentVivo2024] simultaneously. IPANEMAP [saaidiIPANEMAPIntegrativeProbing2020] showed improvements from multi-probe integration but relies on empirical conversions. In our framework, each probe would be modeled with its own parameters while each molecule would have specific $\lambda_{i}$ values, leveraging orthogonal chemical information.
An additional extension concerns RNA-protein interactions that confound probing readouts. Known contacts could be excluded from the loss and constrained against internal pairing; unknown interaction sites could be flagged by localizing discrepancies between predicted and observed profiles.
We note that validation of RNA structure prediction methods typically relies on reference structures from crystallography, cryo-EM, or covariance analysis, each with inherent biases. Crystals often fail to form for conformationally heterogeneous molecules, selecting against dynamic ensembles. Cryo-EM reconstructions typically use only a small fraction of recorded particles, with difficult-to-estimate effects on captured dynamics. Covariance-based methods reflect evolutionarily conserved, biologically functional structures, which may represent only a subset of conformations observed in vitro. These considerations highlight the value of probing-based ensemble methods as a complementary approach to traditional structure determination.
We acknowledge important limitations. The underlying ViennaRNA framework does not account for pseudoknots, critical for many RNAs. More generally, secondary structure alone may not suffice to predict DMS reactivities, which can depend on 3D features [deenalatthaCharacterizing3DRNA2025]. While physical parameters transfer across RNAs under constant conditions, transferability between varying conditions (temperature, buffer, probing duration) is not guaranteed; we recommend fitting similar-condition data together. Broader applicability would require modeling parameter dependence on experimental variables.
1 Data and code availability
The MERGE-RNA source code, together with analysis scripts, is available at https://github.com/giusSacco/MERGE-RNA. Sequencing data for all the experiments performed in this work have been deposited in the European Nucleotide Archive (ENA) under accession number XXXXXXXXX.
2 Competing interests
No competing interest is declared.
3 Author contributions statement
Gi.S., Gu.S., and G.B. conceived and designed the study. Gi.S., Gu.S., and G.B. developed the theoretical modeling and statistical framework, interpreted the data and the results. Gi.S. implemented MERGE-RNA, led the primary data analysis, processed sequencing data, performed model benchmarking, and prepared the figures. G.B. performed exploratory analyses and contributed methodological input. J.L. and R.P.S. designed and performed the DMS-MaP experiments. G.B., Gu.S. and R.P.S. supervised the research. All authors wrote the manuscript, revised it critically, and approved the final version.
4 Acknowledgments
We thank Catherine Eichhorn, Anthony Mustoe and Joseph Yesselman for useful discussions. Anthony Mustoe is also acknowledged for reading a preliminary version of this manuscript and providing useful suggestions. RPS acknowledges the interdisciplinary Thematic Institute IMCBio+, as part of the ITI 2021–2028 program of the University of Strasbourg, CNRS and Inserm, IdEx Unistra (ANR-10-IDEX-0002), SFRI-STRAT’US (ANR 20-SFRI-0012), and EUR IMCBio (ANR-17-EURE-0023) under the framework of the French Investments of the France 2030 Program.
Supplementary Information
Physical Model
The physical model underlying our analysis of RNA chemical probing experiments is illustrated in the left panel of Figure 1 and will be detailed in this section. We derive and justify the expression for the free energy of a secondary structure $s$ in the presence of a chemical probe (as reported in the main text) which enables the prediction of the ensemble of structures, and explain our methodology for producing a prediction of the reactivity profiles at varying probe concentrations, which serve as the theoretical basis for fitting experimental data.
Secondary structure populations
The initial step in our analysis involves modeling the RNA folding into an ensemble of secondary structures, while also accounting for interactions with chemical probes. Our baseline is the prediction of the ViennaRNA package [lorenzViennaRNAPackage202011], specifically employing the RNA.fold_compound.subopt method to obtain an ensemble of suboptimal structures $s$ with their associated free energies $F_{0}(s)$.
The maximum entropy formalism [demartinoIntroductionMaximumEntropy2018] provides a principled framework for integrating experimental data while making minimal adjustments to the baseline predictions from the classical thermodynamic model, ensuring that the resulting distribution of observables gets consistent with experimental measurements. Within this framework, the minimal perturbation to the baseline model that aligns with experimental data is achieved by introducing sequence-specific model parameters $\lambda_{i}$ –one for each nucleotide position– that serve as Lagrange multipliers constraining the ensemble to match experimentally observed properties. These soft constraints can be effectively interpreted as positional corrections of pairing energy, allowing us to capture sequence-dependent effects that may not be fully accounted for in the standard thermodynamic model. This formalism allows to seamlessly integrate experimental data into the thermodynamic framework, and the corrections will be applied only where there are discrepancies between the model and the data, while leaving the areas where there is agreement unchanged. Importantly, the $\lambda_{i}$ parameters are sequence-specific but remain consistent across all concentrations and experimental replicates for the same sequence.
The populations for each secondary structure $s$ is determined by its free energy, which comprises three components: the baseline free energy $E_{0}(s)$, sequence-specific soft constraints $\lambda_{i}$, and the contribution from probe interactions $\Delta F_{\text{[DMS]}}(s)$:
$F(s)$ $=F_{0}(s)+\sum_{i\in\text{paired}(s)}\lambda_{i}+\Delta F_{\text{[DMS]}}(s)$ (7)
The last term explicitly accounts for interactions between chemical probes and RNA structures, which modify the free energy in a concentration-dependent manner, whose explicit form is derived below. The concentration-dependent energy of the probe-RNA interaction is described by chemical potential $\mu$. Note that once a chemical potential for a reference probe concentration is known, the chemical potential for any other concentration can be derived from the ratio of the experimental concentrations as $\mu=\mu_{r}+\mathrm{k_{B}T}\ln\left(\frac{[\text{DMS}]}{[\text{DMS}]_{r}}\right)$, where $\text{DMS}]$ is the concentration of the probe in the experiment and $\text{DMS}]_{r}$ and $\mu_{r}$ are respectively the reference concentration and its chemical potential. Since for most chemical probes the interaction energy depends on local RNA structure (this is especially true for the most widely used SHAPE and DMS), we include an additional parameter $\Delta\mu_{\text{pairing}}$ that represents the energetic penalty for probe interactions with paired nucleotides, such that the effective binding energy for paired nucleotides becomes $\mu^{\prime}\equiv\mu-\Delta\mu_{\text{pairing}}$. This formulation allows us to model the effect of varying probe concentrations with only two parameters ($\mu_{r}$ and $\Delta\mu_{\text{pairing}}$) and thus to combine data from multiple experiments into a single, coherent framework.
Importantly, in what follows we consider only the binding partition function, i.e., the contribution to the total partition function arising from independent probe-RNA binding degrees of freedom at fixed secondary structure $s$. Under the assumption of independent binding across sites, this factorizes as
$Z_{\text{bind}}(s)$ $=\prod_{i:\,s_{i}=0}z_{u}\prod_{i:\,s_{i}=1}z_{p}$ (8)
$=z_{u}^{N_{u}(s)}\,z_{p}^{N_{p}(s)}$ (9)
with single-site partition functions $z_{u}=1+\mathrm{e}^{\beta\mu}$ for unpaired sites and $z_{p}=1+\mathrm{e}^{\beta\mu^{\prime}}$ for paired sites, and where $N_{u}(s)$ and $N_{p}(s)$ denote the numbers of unpaired and paired nucleotides in structure $s$. The corresponding concentration-dependent contribution to the free energy (the last term in Equation 7) is then
$\beta\Delta F_{\text{[DMS]}}(s)$ $=-\ln\big(Z_{\text{bind}}(s)\big)$ (10)
$=-N_{u}(s)\ln z_{u}-N_{p}(s)\ln z_{p}$ (11)
$=-N_{u}(s)\ln(1+\mathrm{e}^{\beta\mu})-N_{p}(s)\ln(1+\mathrm{e}^{\beta\mu^{\prime}})$ (12)
$=-(N-N_{p}(s))\ln(1+\mathrm{e}^{\beta\mu})-N_{p}(s)\ln(1+\mathrm{e}^{\beta\mu^{\prime}})$ (13)
$=-N\ln(1+\mathrm{e}^{\beta\mu})-N_{p}(s)\ln\left(\frac{1+\mathrm{e}^{\beta\mu^{\prime}}}{1+\mathrm{e}^{\beta\mu}}\right)$ (14)
where $N\equiv N_{u}(s)+N_{p}(s)$ is the sequence length. The first term is structure-independent and does not contribute to the populations in the experiment, so it can be safely ignored in our calculations. In practice, this formulation enables us to account for the probe influence on the secondary structure populations by adding $\ln\left(\frac{1+\mathrm{e}^{\beta\mu^{\prime}}}{1+\mathrm{e}^{\beta\mu}}\right)$ as a soft constraint in ViennaRNA for each paired base, allowing us to efficiently compute pairing probabilities within its computational framework. By substituting the expression for $\Delta F_{\text{[DMS]}}(s)$ into Equation 7, we obtain the final expression for the free energy of a secondary structure
$F(s)$ $=\underbrace{F_{0}(s)}_{\text{baseline}}+\underbrace{\sum_{i\in\text{paired}(s)}\lambda_{i}}_{\begin{subarray}{c}\text{site-specific}\\ \text{corrections}\end{subarray}}-\underbrace{\mathrm{k_{B}}TN_{p}(s)\ln\left(\frac{1+\mathrm{e}^{\beta\mu^{\prime}}}{1+\mathrm{e}^{\beta\mu}}\right)}_{\begin{subarray}{c}\text{concentration}\\ \text{dependent perturbation}\end{subarray}}$ (15)
Physical Binding Probability
Once we have outlined how to compute the populations of secondary structures in the presence of a chemical probe, we can derive the expected reactivity profile through the modeling of the chemical probing experiment, beginning from the interaction between the probe and the RNA. The probability of physical binding between the probe and RNA follows a Boltzmann distribution:
$\mathbb{P}(k_{i}=1|s_{i}=0)$ $=\frac{\mathrm{e}^{\beta\mu}}{1+\mathrm{e}^{\beta\mu}}$ (16)
$\mathbb{P}(k_{i}=1|s_{i}=1)$ $=\frac{\mathrm{e}^{\beta\mu^{\prime}}}{1+\mathrm{e}^{\beta\mu^{\prime}}}$ (17)
where $k_{i}\in\{0,1\}$ denotes the physical binding state at position $i$ (1 for probe-bound, 0 for unbound) and $s_{i}\in\{0,1\}$ represents the pairing state (1 for paired, 0 for unpaired).
We make two key assumptions in this model: first, that the probe physical interaction is independent of nucleotide identity, and second, that binding events occur independently at each site, with no interactions between probes. Note that the latter may not hold at high probe concentrations, where cooperative effects and non-linear increases in reactivity can occur, as investigated by [calonaciMolecularDynamicsSimulations2023] and [arnoldCooperativityRNAChemical2023].
Probability of Chemical Modification
The probability of chemical modification at position $i$ depends on both the physical binding event and the subsequent conversion from physical to chemical modification. We assume that when a probe physically interacts with nucleobase $i$, chemical modification occurs with a probability $p_{\text{bind}}(n_{i},s_{i})$, which depends only on the nucleotide identity $n_{i}\in\{\text{A, U, G, C}\}$ and its pairing state $s_{i}$:
$\mathbb{P}(c_{i}=1|s_{i},n_{i})$ $=\mathbb{P}(c_{i}=1|k_{i}=1,s_{i},n_{i})\mathbb{P}(k_{i}=1|s_{i})$ (18)
$\equiv p_{\text{bind}}(n_{i},s_{i})\mathbb{P}(k_{i}=1|s_{i})$ (19)
where $c_{i}\in\{0,1\}$ represents the modification state (1 for chemically modified, 0 for unmodified). While this is the most general formulation and holds for widely used chemical probes such as SHAPE, we note that DMS can react only with unpaired A and C and in a structure independent manner with G and U. We can thus simplify the model and embed the physics of the probe in the model by setting
$p_{\text{bind}}(\text{A},1)$ $=0$
$p_{\text{bind}}(\text{C},1)$ $=0$
$p_{\text{bind}}(\text{G},0)$ $=p_{\text{bind}}(\text{G},1)$
$p_{\text{bind}}(\text{U},0)$ $=p_{\text{bind}}(\text{U},1)$
This leads to a total of four free parameters for the nucleotide-specific binding probabilities, one for each nucleotide type.
The overall probability of chemical modification at site $i$, denoted $\mathbb{P}(c_{i})$, is obtained by averaging over the structural ensemble. Thanks to the assumption that binding and modification events are independent at each site, this average can be computed using the local pairing probabilities $\mathbb{P}(s_{i})$, which are derived from the partition function without enumerating all possible secondary structures. The ensemble-averaged probability of chemical modification is thus given by:
$\mathbb{P}(c_{i})$ $=\left\langle\mathbb{P}(c_{i}|s,n_{i})\right\rangle_{s}$
$=\left\langle\mathbb{P}(c_{i}|s_{i},n_{i})\right\rangle_{s_{i}}$
$=\sum_{s_{i}\in\{0,1\}}\mathbb{P}(s_{i})\mathbb{P}(c_{i}|s_{i},n_{i})$ (20)
Expected Mutation Rate and Systematic Biases
In Mutation Profiling (MaP), chemical modifications induced by probes are detected as mutations in the cDNA during reverse transcription. To accurately model this process, we must establish the relationship between chemical modifications and the observed mutations while accounting for errors introduced during reverse transcription.
Although mutation probabilities could potentially depend on both sequence context and structural environment [verwiltArtifactsBiasesReverse2023], we simplify our model by considering only the dependence on the chemical modification state $c_{i}$, treating other factors as systematic biases (which will be accounted for later in this section). We introduce two parameters $m(c_{i}=1)$ and $m(c_{i}=0)$ that govern respectively mutation probabilities for modified and unmodified nucleotides, according to the following definition:
$\mathbb{P}(\text{mutation at position}~i~|~c_{i})\equiv 1-\mathrm{e}^{-(m(c_{i})+\epsilon_{i})}$ (21)
In this formulation, $m(c_{i}=1)$ and $m(c_{i}=0)$ can be interpreted as sequence-independent probabilities governing respectively false negative and false positive mutations and will be fitted as model parameters. In the following and in the main text we use the shorthand notation $m_{1}\equiv m(c_{i}=1)$ and $m_{0}\equiv m(c_{i}=0)$. The position-specific $\epsilon_{i}$ capture systematic biases common across all experimental replicates and probe concentrations for the same RNA sequence. These $\epsilon_{i}$ values are calibrated to ensure that the model-predicted reactivity profile at zero probe concentration aligns with the control experiment, effectively serving as site-specific offsets common to all experimental conditions.
To compute the model prediction of the experimental reactivity profile, we average over all possible modification states, weighted by their probabilities:
$M_{i}$ $\equiv\langle M(c_{i})\rangle_{c_{i}}$ (22)
$=\sum_{j\in\{0,1\}}\mathbb{P}(c_{i}=j)M(c_{i}=j)$ (23)
$=\sum_{j\in\{0,1\}}\mathbb{P}(c_{i}=j)(1-\mathrm{e}^{-(m(j)+\epsilon_{i})})$ (24)
$=1-\mathrm{e}^{-\epsilon_{i}}\sum_{j\in\{0,1\}}\mathbb{P}(c_{i}=j)\mathrm{e}^{-m_{j}}$ (25)
$=1-\mathrm{e}^{-\epsilon_{i}}\left[(1-\mathbb{P}(c_{i}=1))\mathrm{e}^{-m_{0}}+\mathbb{P}(c_{i}=1)\mathrm{e}^{-m_{1}}\right]$ (26)
$=1-\mathrm{e}^{-\epsilon_{i}}\left[\mathrm{e}^{-m_{0}}+\mathbb{P}(c_{i}=1)(\mathrm{e}^{-m_{1}}-\mathrm{e}^{-m_{0}})\right]$ (27)
This equation establishes a critical link between experimentally observed mutation rates and the underlying RNA structural ensemble through the chemical modification probabilities $\mathbb{P}(c_{i}=1)$, which themselves depend on the pairing state probabilities $\mathbb{P}(s_{i})$ derived earlier.
Notably, the first term in the square brackets, containing $m_{0}$, produces a common background signal. When control experiments are available, this term can be incorporated into the systematic biases $\epsilon_{i}$; otherwise, it must be fitted as a constant background term.
Model Parameters Summary
Before moving on to how to obtain the optimal model parameters from experimental data, we summarize here all the parameters of our model that require fitting:
  • Physical parameters, which describe the physics of the chemical probing experiment. These are shared across all RNA sequences and probe concentrations.
    • $\mu_{r}$: quantifies the energy of probe-RNA interaction at a reference probe concentration, from which we derive chemical potentials for all other concentrations.
    • $\Delta\mu_{\text{pairing}}$: represents the energetic penalty for probe interaction with paired nucleobases.
    • $p_{\text{bind}}(n_{i},s_{i})$: is the probability of chemical modification given nucleotide identity $n_{i}$ and pairing state $s_{i}$. These are four for DMS, one for each nucleotide type.
    • $m_{1}$ and $m_{0}$: are parameters governing respectively false negative and false positive mutation probabilities during the cDNA transcription. If control experiments are not available, $m_{0}$ gets absorbed into the systematic biases $\epsilon_{i}$ and does not need to be fitted.
  • Soft constraints, which are sequence-specific but consistent across all experimental conditions for a given RNA:
    • $\lambda_{i}$: are site-specific soft constraints modifying the pairing energy of the baseline thermodynamic model.
Worth mentioning in this summary–although not fitted–are the $\epsilon_{i}$, which are calibrated from the control experiment and incorporate site-specific systematic errors shared across replicates and concentrations, and effectively serve as offsets to the predicted mutation profiles. They enable the decoupling of experimental biases from the thermodynamic and concentration-dependent effects, allowing the model to focus on physically relevant signal.
Model Training
Loss function
To fit the model parameters to experimental data, we define a loss function that quantifies the discrepancy between the expected reactivity profiles predicted by our model (as derived in the previous section) and the observed mutation rates from chemical probing experiments.
The binomial nature of the mutation counts, which arise from a series of independent Bernoulli trials (each read can either show a mutation or not), motivates the choice of a binomial likelihood for our loss function. Within this framework, the probability of observing $\mathcal{M}_{i}$ mutations out of a coverage of $n_{i}$ reads under predicted mutation rate $M_{i}$ is given by:
$\mathbb{P}(\mathcal{M}_{i})$ $=\text{Binomial}(\mathcal{M}_{i},n_{i},M_{i})$ (28)
$=\begin{pmatrix}n_{i}\\ \mathcal{M}_{i}\end{pmatrix}(1-M_{i})^{n_{i}-\mathcal{M}_{i}}$ (29)
where $i$ runs over all the data points, i.e. sequences, positions, concentrations and replicates. We thus define the loss function as the negative log-likelihood of observing the experimental mutation counts $\mathcal{M}_{i}^{\text{exp}}$ given the model.
$\mathcal{L}$ $=\sum_{i}-\ln\mathbb{P}(\mathcal{M}_{i}^{\text{exp}}=\mathcal{M}_{i})$ (30)
$=\sum_{i}-\ln\left[\begin{pmatrix}n_{i}\\ \mathcal{M}_{i}^{\text{exp}}\end{pmatrix}M_{i}^{\mathcal{M}_{i}^{\text{exp}}}(1-M_{i})^{n_{i}-\mathcal{M}_{i}^{\text{exp}}}\right]$ (31)
Fit of parameters
The model fitting process involves optimizing a set of parameters to best reproduce experimental data. We implemented the optimization procedure in a custom Python script utilizing the scipy.optimize.minimize function [virtanenSciPy10Fundamental2020] with the L-BFGS-B algorithm [byrdLimitedMemoryAlgorithm1995]. This limited-memory quasi-Newton method is particularly well-suited for high-dimensional optimization problems, as it efficiently approximates the inverse Hessian matrix while maintaining reasonable memory requirements. The algorithm’s ability to handle bound constraints is especially valuable for ensuring physically meaningful parameter values throughout the optimization process. Specifically, we constrained the parameters to physically plausible ranges: the chemical potential at reference concentration, $\mu_{r}$, was bounded between -5 and 5 kcal/mol; the energetic penalty for pairing, $\Delta\mu_{\text{pairing}}$, was constrained to be non-negative, with an upper bound of 10 kcal/mol; the chemical modification probabilities, $p_{\text{bind}}$, were constrained to the range $[0,1]$, as is required for probabilities; the false positive mutation parameter, $m_{0}$, was kept small within $[0,0.1]$; and the false negative parameter, $m_{1}$, was bounded within $[0.1,10]$.
The optimization is performed in a two-stage process to minimize the loss function. In the first stage, we fit the physical parameters that are common to all experiments, keeping the sequence-specific soft constraints ($\lambda_{i}$) fixed. This global optimization starts from random initializations and runs until convergence. In the second stage, the optimized physical parameters are held constant while we fit the $\lambda_{i}$ soft constraints independently for each RNA sequence. This hierarchical strategy allows us to first establish a robust physical model and then refine the secondary structure populations with sequence-specific adjustments. During the optimization, the $\lambda_{i}$ parameters are constrained to the range [1,1][-1,1] kcal/mol to ensure proper regularization and avoid overfitting.
Experimental Procedures
In vitro transcription (IVT)
A single-stranded full-length DNA oligonucleotide (5’-ATGGTCTGCTGGAGGTTGGAATAGTTGTGAGTTGAAGTGGGGATTTTTTAATCCCCACTTCAACTCACAACTATTCCAATTTTTTGGAATAGTTGTGAGTTGAAGTGGGGATTGGCTGTGGCATAATG-3’) was chemically synthesized (Integrated DNA Technologies). This full-length sequence was amplified by PCR using primers carrying a T7 promoter. The primers used were Full-T7-RV (5’-TAATACGACTCACTATAGGGCATTATGCCACAGCCAATCCCCACTTC-3’) and Full-FW (5’-ATGGTCTGCTGGAGGTTGGAATAGTTGTGAG-3’).
[…]
RNA refolding
DNase-treated IVT products were column-purified and used for RNA refolding. For the initial refolding step, RNA was diluted in ultrapure water to an RNA concentration of 26 nM, and refolding buffer was added to reach the following final concentrations: EDTA (0.5 mM), HEPES (200 mM, pH 7.5), NaCl (300 mM). The mixture was heated at 95 C for 1.5 min to denature RNA secondary structures and then immediately placed on ice. While on ice, MgCl2 was added to give a final concentration of 5 mM and an RNA concentration of 23.5 nM. Samples were then incubated at 37 C for 30 min to allow RNA refolding. Refolded RNA was stored at 80-80 C until further use.
DMS probing
For dimethyl sulfate (DMS) probing, reactions were prepared with refolded RNA at a final concentration of 23.5 nM. DMS (Sigma–Aldrich) stock solutions were prepared in 100% ethanol such that, upon addition to the RNA samples, the final DMS concentrations were 20, 40, 60, 80, or 100 mM. $\beta$-mercaptoethanol (Sigma–Aldrich) was diluted in PBS such that, upon addition to the reaction, its final concentration was 1000 mM. For each DMS concentration, one negative control (receiving an equal volume of 100% ethanol instead of DMS) and two technical replicates with DMS were prepared. Refolded RNA samples were kept on ice, mixed thoroughly with the appropriate DMS working solution or with 100% ethanol for negative controls, and immediately transferred to an incubator pre-equilibrated to 37 C for 7 min. Reactions were then placed back on ice, and $\beta$-mercaptoethanol working solution was added to each reaction and mixed thoroughly to reach its final concentration of 1000 mM and to quench DMS modification.
Reverse transcription
[…]
PCR amplification
[…]
Nanopore sequencing
[…]
Mapping and counting of mutations
Sequencing reads were processed with an RNAFramework-based [incarnatoRNAFrameworkAllinone2018] workflow using the functions rf-map and rf-count, following a similar pipeline to [morandiGenomescaleDeconvolutionRNA2021] with minor adaptations for our constructs. For each library, we generated a Bowtie2 [langmeadFastGappedreadAlignment2012] index from the reference sequence using bowtie2-build and produced the companion FASTA index with samtools faidx [danecekTwelveYearsSAMtools2021]. Reads were aligned to the indexed reference with rf-map -cq5 20 -cqo -mp ’--very-sensitive-local’ -b2 -bi ../{ref_fasta}_index, where {ref_fasta}_index are the files produced by bowtie2-build. The resulting BAM files were coordinate sorted and indexed (samtools sort/index) before quantification.
Mutation profiles were obtained with rf-count [incarnatoRNAFrameworkAllinone2018] using the options recommended by Ref.[morandiGenomescaleDeconvolutionRNA2021] (-m -ds 75 -na -ni -md 3).These parameters report mutational load rather than RT stops (-m), discard reads shorter than 75 nt (-ds 75), ignore ambiguous deletions (-na) and indels (-ni), and cap deletions at three nucleotides (-md 3). The full command sequence executed for each FASTQ file is summarized below:
# build indices of reference
bowtie2-build --quiet {ref_fasta} {ref_fasta}_index
samtools faidx {ref_fasta}
# alignment (see above)
rf-map \
    -cq5 20 \
    -cqo \
    -mp ’--very-sensitive-local’ \
    -b2 \
    -bi {ref_fasta}_index \
[…]
    {fastq_file}
# sorting and indexing the BAM file
samtools sort {bam_file} -o {sorted_bam_file}
samtools index {sorted_bam_file}
# count mutations to obtain mutation profile and mutation map
rf-count \
    -f {ref_fasta} \
    -m \
    -ds 75 \
    -na \
    -ni \
    -md 3 \
    {sorted_bam_file}
For the reference structured RNAs (HCV IRES, bacterial RNase P, V. cholerae glycine riboswitch, Tetrahymena ribozyme, HC16), we instead adopted the filtering scheme described by the original work [bohnNanoDMSMaPAllowsIsoformspecific2023].
rf-count -r -p {num_threads} \
    -mf {primer_mask} \
    -f {reference_fasta} \
    --only-mut ’G>Y;A>B;C>D;T>V’ \
    -m \
    -nd \
    -ni \
    -q {minq} \
    -eq 10
All downstream structural analyses reported in the main text rely on the coverage and mutation rates generated by this standardized workflow.
Supplementary Figures and Tables
Training sets $\mu_{r}$(kcal/mol) $\Delta\mu_{\text{pairing}}$ (kcal/mol) $p_{\text{bind}}(\text{A},\text{unpaired})$ $p_{\text{bind}}(\text{C},\text{unpaired})$ $p_{\text{bind}}(\text{G})$ $p_{\text{bind}}(\text{U})$ $m_{1}$
hc16 + HCV + V.chol. -0. 08 0. 34 1. 00 0. 60 0. 48 0. 05 0. 53
hc16 + RNaseP + HCV -0. 05 0. 49 1. 00 0. 57 0. 57 0. 05 0. 50
hc16 + RNaseP + Tetrah. -0. 02 0. 15 0. 89 0. 56 0. 28 0. 02 0. 56
hc16 + RNaseP + V.chol. -0. 36 0. 03 1. 00 0. 66 0. 31 0. 04 1. 05
hc16 + Tetrah. + HCV -0. 06 0. 61 0. 98 0. 44 0. 58 0. 03 0. 50
hc16 + Tetrah. + V.chol. -0. 07 0. 40 0. 91 0. 64 0. 40 0. 03 0. 53
RNaseP + HCV + V.chol. 0. 08 0. 18 0. 99 0. 58 0. 35 0. 05 0. 39
RNaseP + Tetrah. + HCV -0. 08 0. 24 0. 97 0. 51 0. 35 0. 04 0. 53
RNaseP + Tetrah. + V.chol. -0. 05 0. 79 0. 91 0. 61 0. 41 0. 07 0. 54
Tetrah. + HCV + V.chol. -0. 09 0. 50 1. 00 0. 49 0. 57 0. 05 0. 45
Table S1: Physical parameters obtained by training on different triplets of RNA systems.
Figure S1:
Figure S2:
Figure S3:
Figure S4:
Figure S5:
Figure S6:
Figure S7:
Figure S8:
Figure S9:
Figure S10:
Figure S11:
Figure S12:
Table S2: Loop co-occupancy inferred from baseline thermodynamic ensembles and after MERGE-RNA refinement. Entries report the percentage of structures in which Loop1 and Loop2 are formed.
Baseline MERGE-RNA
Turner
Loop1
Loop2 0 1
0 0% 49%
1 51% 0%
Loop1
Loop2 0 1
0 0% 27%
1 16% 57%
Andr.
Loop1
Loop2 0 1
0 0% 19%
1 81% 0%
Loop1
Loop2 0 1
0 0% 42%
1 16% 42%